# HGRN2: Gated Linear RNNs with State Expansion

### v1

@@ Abstract
Hierarchically gated linear RNN (HGRN, Qin et al. 2023c) has demonstrated competitive training speed and performance in language modeling, while offering efficient inference. However, the recurrent state size of HGRN remains relatively small, which limits its expressiveness. To address this issue, inspired by linear attention, we introduce a simple outer-product-based state expansion mechanism so that the recurrent state size could be significantly enlarged without introducing any additional parameters. The linear attention form also allows for hardware-efficient training. Our extensive experiments verify the advantage of HGRN2 over HGRN1 in language modeling, image classification, and Long Range Arena. Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
@@ 1 Introduction
Large language models (LLMs) have achieved great empirical success in recent years. However, serving Transformer-based LLMs is expensive due to the expensive KV cache management. Recurrent neural networks (RNNs), on the other hand, offer linear inference complexity with constant state size, which is ideal for serving. Therefore, there is great interest in studying parallelizable linear recurrent models such as linear RNNs (Peng et al., 2023; Orvieto et al., 2023; Qin et al., 2023c; De et al., 2024), linear attention (Sun et al., 2023b; Qin et al., 2023b; Yang et al., 2023; Arora et al., 2024; Peng et al., 2024), and state space models (Gu et al., 2022a; Smith et al., 2023; Gu & Dao, 2023).
@@
RNNs have a fixed recurrent state size to encode all historical information. Therefore, it is important for RNNs (i) to better exploit the fix-sized states and (ii) to increase the recurrent state size. Recent improvements in linear RNN are exactly on this track, including the use of data-dependent decays and state expansion techniques.
@@
Data-dependent decays (also known as forget gates) are of great importance for RNNs (van der Westhuizen & Lasenby, 2018), which allows them to selectively retain useful information while erasing irrelevant information, so that the fixed-size recurrent state could be better exploited. HGRN (Qin et al., 2023c) first highlighted the importance of data-dependent decays for linear RNNs. Since then, many recent linear recurrent models have also employed data-dependent decays, such as Mamba (Gu & Dao, 2023), Gated Linear Attention (GLA, Yang et al. 2023), Griffin (De et al., 2024), and RWKV-6 (Peng et al., 2024).
@@
However, HGRN did not increase the recurrent state size size, greatly restricted by the limited memory capacity, which prevents it from achieving LLaMa-like (Touvron et al., 2023a; b) language modeling performance, as noted in Qin et al. (2024). Recent state-of-the-art linear recurrent models, such as Mamba, GLA, and RWKV-6, additionally use state expansion techniques, which greatly increase the recurrent state size and thereby enhance the memory capacity.
@@
In this work, we propose HGRN2, with the aim of increasing the size of the recurrent state for HGRN while retaining both parameter efficiency and training efficiency. We first explore structured matrices to directly expand the state size in a parameter-efficient manner. Empirically, we found that they improved the language modeling performance, but still faced with the training inefficiency issue, thus they cannot scale the recurrent state size well. Inspired by linear attention, we then explore using a nonparameteric outer product-based state expansion mechanism, which allows efficient scaling of the size of the recurrent state during training without introducing any additional parameters. Thanks to the gated linear attention form, we could borrow the hardware-efficient GLA training algorithm described in Yang et al. (2023) for large-scale experiments.
We extensively evaluate HGRN2 in language modeling, image classification, and the Long Range Arena benchmark. HGRN2 outperforms HGRN1 in all these benchmarks, showing the significant benefit of using state expansion. Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
@@ 2.1 Gated linear RNN
Given input $\mathbf{x}\in\mathbb{R}^{N\times d}$ where the sequence length is $N$ and the model dimension is $d$. A minimalist gated linear recurrent layer transforms the input $\mathbf{x}$ into hidden states $\mathbf{h}\in\mathbb{R}^{N\times d}$ and the output $\mathbf{y}\in\mathbb{R}^{N\times d}$, as defined below.
@@
$\mathbf{i}_{t}$ $=\mathbf{V}\mathbf{x}_{t}+\mathbf{b}_{v},$
$\mathbf{o}_{t}$ $=\tau\left(\mathbf{W}\mathbf{x}_{t}+\mathbf{b}_{w}\right),$
@@
where $\odot$ is element-wise product; $\sigma$ is the sigmoid function and $\tau$ is nonlinear activation (we choose to use $\mathrm{SiLU}$); $\mathbf{i}_{t}$ is the input vector; $\mathbf{g}_{t},\mathbf{o}_{t}$ are the forget gate and output gate, respectively. The input gate is tied to the forget gate: $1-\mathbf{g}_{t}$, which is commonly used in many gated RNNs such as GRU (Chung et al., 2014).
@@
Compared to Eq. 1, HGRN makes two adjustments: (i) complex-valued recurrence. (ii) forget gates with monotonically increased lower bound values from bottom layers to upper layers.
For (i), similarly to the findings of Gu & Dao (2023) and De et al. (2024), we empirically found that the complex-valued recurrence is not necessary, as shown in Table 1. We speculate that the reason Qin et al. (2023c) found it useful is due to state expansion: The complex-valued recurrent state is twice the size of that in the real-valued recurrent state. If we directly expand the real-valued recurrent state size from $d$ to $2d$, the language modeling performance on Wikitext-103 corpus is even better. Therefore, we only consider the real-valued recurrence thereafter.
@@
Real HGRN1 $2d$ 23.81 24.55 47.82
For (ii), suppose that the total number of layers is $L$, and HGRN introduces a data-independent learnable matrix $\Gamma\in\mathbb{R}^{L\times d}$. $\Gamma_{i}$ will be the lowest values of the forget gate for the $i$th layer in all time steps. HGRN argues that this lower bound should be monotonically increasing from bottom to top, encouraging the bottom layers to model short-term local dependencies and the upper layers to model long-term dependencies. To enforce such monotonicity, HGRN used the cumulative softmax operator cumax [^1] This was first introduced by Shen et al. (2018):
$\beta:=\texttt{cumax}(\Gamma)=\texttt{cumsum}(\texttt{softmax}(\Gamma,\mathrm{dim=0}),\mathrm{dim=0})\in\mathbb{R}^{L\times d},\beta^{i}=[\beta]_{i}\in\mathbb{R}^{d}.$
To avoid the lower bound being one in the highest layer, HGRN subtracts all $\beta$ by $\beta^{0}$, so that the lower bound for the first layer is zero. After obtaining the lower bound values, the old forget gate $\mathbf{g}_{t}$ learns the residuals instead, resulting the new forget gate $\mathbf{f}_{t}$
@@
where we use superscript to indicate the layer index.
Figure 1: The neural architecture of HGRN2. Each HGRN2 layer includes a token mixer layer HGRU2 and a channel mixerlayer GLU. HGRU2 employs recurrent computation through Eq. 3, where $\mathbf{i}_{t}$ is the input vector, $\mathbf{g}_{t}$ is the forget gate (not lower bounded), $\beta^{i}$ is the lower bound of the forget gate value, $\mathbf{o}_{t}$ is the output gate for layer $i$.
@@ 3.1 Explorations of state expansion methods
The goal is to scale the size of the HGRN recurrent state from $d$ to $nd$, where $n$ is the expansion ratio of the state. However, if we use the original parameterization in Eq. 1, $\mathbf{U},\mathbf{V},\mathbf{W}$ will be of size $d\times nd$, which is very parameter inefficient when $n$ is large. Ideally, the number of parameters should be around $d^{2}$ as in the original case for each projection. First, we consider the following structured matrices, including low-rank, group linear transformation (GLT, Mehta et al. 2018; 2021), group linear transformation with group interaction (GLTI), Khatri-Rao product (KRP) and Kronecker product (KP), to replace the dense projection matrix $\mathbb{R}^{d}\rightarrow\mathbb{R}^{nd}$ as described in Table 2.
@@
Table 2: We use Einstein Summation notation to represent Parameter Efficient State Expand (PESE) methods. We use Blue to represent the input, Black to represent data independent weight and Red to represent the output.
PESE Naive LR GLT
Equation ${{\color[rgb]{0,0,1}d},d\ nd\to{\color[rgb]{1,0,0}nd}}$ ${\color[rgb]{0,0,1}d},d\ r,r\ nd\to{\color[rgb]{1,0,0}nd}$ $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n\ e},n\ e\ d\to{\color[rgb]{1,0,0}(n\ d)}$
Prameters $nd^{2}$ $dr(n+1)\approx d^{2}$ $d^{2}$
PESE GLTI KRP KP
Equation $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n^{\prime}\ e},n^{\prime}\ e\ d,n^{\prime}\ n\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},n\ d\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},d\ d,n\to{\color[rgb]{1,0,0}(n\ d)}$
Parameters $d^{2}+n^{2}$ $nd$ $d^{2}+n$
@@
After obtaining expanded $\mathbf{g},\mathbf{i},\mathbf{o}$, we feed them into element-wise gated linear recurrent layers as in Eq 1 and  2, resulting in the output vector $\mathbf{y}_{t}\in\mathbb{R}^{n\times d}$. To project the expanded dimension back to the original dimension, we simply marginalize the dimension corresponding to $n$.
@@
The result is shown in Table 3. We found that state expansion generally improves the performance and that the low-rank matrix performs the best among these candidates.
@@
However, these PESE methods face the training inefficiency issue, as they need to conduct element-wise linear recurrence in high dimension (i.e. $nd$). Since these element-wise operations cannot leverage tensor cores (a fast matrix multiply unit on GPUs), the dramatically increasing FLOPs and I/O costs would significantly slow down the training when $n$ is large. We notice that this is similar to the case in Mamba [^2] Though Mamba has an attention view (Ali et al., 2024) similar to that in linear attention, the attention computation cannot be written as matrix multiply like linear attention, and thus does not facilitate tensor core-based GPU acceleration., so Mamba needs a relatively small expansion ratio (i.e., $n=16$) and also a custom I/O efficient CUDA implementation to achieve a reasonably fast running speed.
@@
In the next subsection, we explore another strategy that does not replace the dense projection matrices with structured ones, but changes the element-wise gating operations in Eq.1 to some other matrix/vector operations similar to linear attention, which allows for efficient training.
Table 3: PESE Ablation. We also provide the loss of Transformer (Xfmr), LLaMA (Xfmr++), and HGRN1 for references. $n$ standards for Expand Ratio. All experiments were trained on 10 billion tokens in a subset of the Pile.
Method $n$ Loss Params (B)
Xfmr - 2.362 0.380
Xfmr++ - 2.201 0.386
HGRN1 1 2.343 0.379
LR 4 2.246 0.385
8 2.247 0.386
GLT 4 2.333 0.386
GLTI 4 2.267 0.386
KRP 4 2.337 0.386
KP 4 2.332 0.386
HGRN2 4 2.255 0.385
8 2.237 0.385
128 2.195 0.385
@@ 3.2 HGRN2
The modification from HGRN1 to HGRN2 is simple yet effective. For the input gating mechanism, HGRN2 replaces the element-wise product with the outer product for state expansion. Consequently, $\mathbf{h}_{t}\in\mathbb{R}^{d\times d}$ and HGRN2 first diagonalize the forget gate vector and use the matrix dot product to update the hidden state. For the output gating mechanism, HGRN2 replaces the element-wise product with matrix-vector multiplication to project the expanded state to the original dimension size. The recurrent equation of HGRN2 is as follows,
$\mathbf{h}_{t}$ $=\mathrm{Diag}\{\mathbf{f}_{t}\}\cdot\mathbf{h}_{t-1}+(1-\mathbf{f}_{t})\otimes\mathbf{i}_{t}\in\mathbb{R}^{d\times d},$ (3)
$\mathbf{y_{t}}$ $=\mathbf{o}_{t}\cdot\mathbf{h}_{t}\in\mathbb{R}^{1\times d},$
where $\mathrm{Diag}$ stands for diagonalizing vectors, $\cdot$ is the matrix dot product, $\otimes$ is outer product. Note that this recurrent computation resembles that in GLA and RWKV-5/6.
@@
Note that the complexity of recurrence increases dramatically from $O(BNd)$ to $O(BNd^{2})$ due to state expansion. Therefore, we introduce multi-head HGRN (similar to that in linear attention) such that the complexity is reduced to $O(BNd^{2}/H)$ for the number of heads $H$ and the state size effectively becomes $d^{2}/H$, i.e., the expansion ratio $n=d_{h}=d/H$ [^3] See Bolya et al. (2022) for more detailed complexity analysis. We conduct an ablation study on the expand ratio(head dimension) $n=\frac{d}{H}$ as shown in Figure 2. We can see that state expansion is very effective in improving language modeling performance. However, when the head dimension (i.e., state expansion ratio) is larger than 128, the performance gain is not too high. To balance computational cost and performance, we chose to use $d_{h}=128$ for the main experiments.
Figure 2: Expand Ratio (head dimension) Ablation. We tested the relationship between PPL (Perplexity) and expand ratio on the Wikitext-103 (Merity et al., 2017)(left) and a subset of the Pile (Gao et al., 2020) dataset(right). It can be seen that as the expand ratio increases, the PPL decreases.
@@
Also note that this form of recurrence is similar to that of Gated Linear Attention (Yang et al., 2023) except for concrete parameterization. We list the correspondence between two parameterizations in Table 4. We can see that the output gate in HGRN2 amounts to the query in GLA, while the output gate in GLA is removed in HGRN2. The key vector in GLA corresponds to the input gate in HGRN2, which is tied to the forget gate, thus saving parameters.
@@
-- $\mathbf{o}$ (output gate)
Table 4: The correspondence between HGRN2 and GLA. Compared to GLA, HGRN2 removes GLA’s output gate and ties key vector to the forget gate (in terms of GLA terminology), thus saving parameters.
@@
Thanks to the similar computation structure to GLA, we can directly leverage their chunkwise algorithm and their highly optimized kernels  [^4] https://github.com/sustcsonglin/flash-linear-attention for hardware-efficient large-scale training. We refer the reader to their paper for more details.
Ending remarks.
Although HGRN2 shares many similarities with GLA, we believe that HGRN2 offers a unique perspective that is different from linear attention, starting from the approach of gated linear RNN. For example, it is not immediately apparent from the perspective of linear attention why key vectors should be constrained within the range of (0, 1) such as in Schlag et al. (2021) ; and why key vector and forget gate value should sum up to one. It becomes quite intuitive when one starts with the gated linear RNN framework and aims to explore state expansion.

### v2

@@ Abstract
Hierarchically gated linear RNN (HGRN, Qin et al. 2023c) has demonstrated competitive training speed and performance in language modeling while offering efficient inference. However, the recurrent state size of HGRN remains relatively small, limiting its expressiveness. To address this issue, we introduce a simple outer product-based state expansion mechanism, which significantly enlarges the recurrent state size without introducing any additional parameters. This enhancement also provides a linear attention interpretation for HGRN2, enabling hardware-efficient training. Our extensive experiments verify the advantage of HGRN2 over HGRN consistently across different settings and competitive with other recurrent models.
@@ 1 Introduction
Large language models (LLMs) have achieved significant empirical success in recent years. However, serving Transformer-based LLMs is costly due to the expensive KV cache management. Recurrent neural networks (RNNs), on the other hand, offer linear inference complexity with constant state size, making them ideal for serving. Consequently, there is substantial interest in studying parallelizable linear recurrent models, such as linear RNNs (Peng et al., 2023; Orvieto et al., 2023; Qin et al., 2023c; De et al., 2024), linear attention (Sun et al., 2023; Qin et al., 2023b; Yang et al., 2023; 2024; Arora et al., 2024), and state space models (Gu et al., 2022a; Smith et al., 2023; Gu & Dao, 2023; Dao & Gu, 2024).
@@
RNNs have a fixed recurrent state size to encode all historical information. Therefore, it is important for RNNs to (i) utilize the fixed-sized states effectively and (ii) increase the recurrent state size to enhance memory capacity. Recent improvements in linear RNNs follow this approach, incorporating techniques such as data-dependent decays and state expansion.
@@
Data-dependent decays (also known as forget gates) are crucial for RNNs (van der Westhuizen & Lasenby, 2018), allowing them to selectively retain useful information while erasing irrelevant information. This enables the fixed-size recurrent state to store only important information more efficiently. HGRN (Qin et al., 2023c) first emphasized the importance of data-dependent decays for linear RNNs. Many recent linear recurrent models, such as Mamba (Gu & Dao, 2023), Gated Linear Attention (GLA, Yang et al. 2023), Griffin (De et al., 2024), and RWKV-6 (Peng et al., 2024), also employ data-dependent decays.
@@
However, HGRN did not increase the recurrent state size, which is greatly restricted by limited memory capacity. This limitation prevents it from achieving LLaMa-like (Touvron et al., 2023a; b) language modeling performance, as noted in Qin et al. (2024). Recent state-of-the-art linear recurrent models, such as Mamba, GLA, and RWKV-6, have addressed this issue by employing state-expansion techniques. These techniques significantly increase the recurrent state size and thereby enhance memory capacity, which has been shown to be crucial for language modeling performance and directly correlated with retrieval ability (Arora et al., 2024).
@@
In this work, we propose HGRN2, which aims to increase the recurrent state size for HGRN while retaining both parameter and training efficiency. We first explore structured matrices to expand the state size directly in a parameter-efficient manner. Empirically, we found that this approach improves language modeling performance but still encounters training inefficiencies, which limit the scaling of the recurrent state size. Inspired by linear attention, we then explore using a non-parametric outer product-based state expansion mechanism. This approach allows for efficient scaling of the recurrent state size during training without introducing additional parameters. Due to the matrix multiply form of linear attention, we can leverage the hardware-efficient linear attention training algorithm described in Yang et al. (2023); Qin et al. (2024) for large-scale experiments. As a result, HGRN2 can be regarded as an improved parameterization of GLA.
We extensively evaluate HGRN2 across various tasks, demonstrating that it consistently outperforms HGRN1 in multiple domains. In language modeling, we show HGRN2 to be highly competitive compared to other subquadratic efficient models.
@@ 2.1 Gated linear RNN
Given input $\mathbf{x}\in\mathbb{R}^{N\times d}$, where the sequence length is $N$ and the model dimension is $d$, a minimalist gated linear recurrent layer (Martin & Cundy, 2018) transforms the input $\mathbf{x}$ into hidden states $\mathbf{h}\in\mathbb{R}^{N\times d}$ and the output $\mathbf{y}\in\mathbb{R}^{N\times d}$, as defined below:
@@
$\mathbf{i}_{t}$ $=\tau\left(\mathbf{V}\mathbf{x}_{t}+\mathbf{b}_{v}\right),$
$\mathbf{o}_{t}$ $=\sigma\left(\mathbf{W}\mathbf{x}_{t}+\mathbf{b}_{w}\right),$
@@
where $\odot$ denotes element-wise product; $\sigma$ is the sigmoid function, and $\tau$ is a nonlinear activation function (we choose to use $\mathrm{SiLU}$); $\mathbf{i}_{t}$ is the input vector; $\mathbf{g}_{t}$ and $\mathbf{o}_{t}$ are the forget gate and output gate, respectively. The input gate is tied to the forget gate as $1-\mathbf{g}_{t}$, a common approach used in many gated RNNs such as GRU (Chung et al., 2014).
@@
Compared to Eq. 1, HGRN makes two adjustments: (i) complex-valued recurrence and (ii) forget gates with monotonically increased lower bound values from bottom layers to upper layers.
For (i), similar to the findings in Gu & Dao (2023) and De et al. (2024), we empirically found that complex-valued recurrence is not necessary, as shown in Table 1. The reason why HGRN found it useful is due to state expansion: the complex-valued recurrent state is twice the size of that in the real-valued recurrent state. If we directly expand the real-valued recurrent state size from $d$ to $2d$, the language modeling performance on the Wikitext-103 corpus is even better. Therefore, we only consider the real-valued recurrence thereafter.
@@
Real HGRN1 $2d$ 24.04 24.64 45.46
For (ii), suppose the total number of layers is $L$. HGRN introduces a data-independent learnable matrix $\Gamma\in\mathbb{R}^{L\times d}$, where $\Gamma_{i}$ represents the lowest values of the forget gate for the $i$-th layer at all time steps. HGRN argues that this lower bound should be monotonically increasing from bottom to top, encouraging the bottom layers to model short-term local dependencies and the upper layers to model long-term dependencies. To enforce this monotonicity, HGRN uses the cumulative softmax operator cumax (Shen et al., 2018):
$\beta:=\texttt{cumax}(\Gamma)=\texttt{cumsum}(\texttt{softmax}(\Gamma,\text{dim}=0),\text{dim}=0)\in\mathbb{R}^{L\times d},\quad\beta^{i}=[\beta]_{i}\in\mathbb{R}^{d}.$
To prevent the lower bound from reaching one in the highest layer, HGRN subtracts all $\beta$ values by $\beta^{0}$, making the lower bound for the first layer zero. After obtaining the lower bound values, the forget gate $\mathbf{g}_{t}$ learns residuals instead, resulting in a new forget gate $\mathbf{f}_{t}$:
@@
where the superscript indicates the layer index. This additive lower bound approach has been shown to mitigate the issue of saturated gates (Gu et al., 2020).
@@ 3.1 Explorations of state expansion methods
The goal of this work is to scale the size of the HGRN recurrent state from $d$ to $nd$, where $n$ is the state expansion ratio. However, if we use the original parameterization in Eq. 1, the matrices $\mathbf{U},\mathbf{V},\mathbf{W}$ will have dimensions $d\times nd$, which becomes very parameter inefficient when $n$ is large. Ideally, the number of parameters should be around $d^{2}$, as in the original case for each projection. To achieve this, we first consider using structured matrices (e.g., low-rank matrices) to replace the dense projection matrix $\mathbb{R}^{d}\rightarrow\mathbb{R}^{nd}$, as described in Table 2.
@@
Table 2: Parameter Efficient State Expansion (PESE) methods using Einstein Summation notation. Blue represents the input, Black represents data-independent weights, and Red represents the output. We list the Einstein Summation for low-rank (LR), group linear transformation (GLT), group linear transformation with interaction (GLTI), Khatri-Rao product (KRP), and Kronecker product (KP).
      Method       Equation       Parameter #
      Naive       ${{\color[rgb]{0,0,1}d},\mathbf{d\ nd}\rightarrow{\color[rgb]{1,0,0}nd}}$       $nd^{2}$
      LR       ${\color[rgb]{0,0,1}d},\mathbf{d\ r,\ r\ nd}\rightarrow{\color[rgb]{1,0,0}nd}$       $dr(n+1)\approx d^{2}$
      GLT              $\color[rgb]{0,0,1}d=(n\ e)\rightarrow n\ e$       ${\color[rgb]{0,0,1}n\ e},\mathbf{n\ e\ d}\rightarrow{\color[rgb]{1,0,0}n\ d}$       $d^{2}$
      GLTI              $\color[rgb]{0,0,1}d=(n\ e)\rightarrow n\ e$       ${\color[rgb]{0,0,1}n\ e},\mathbf{n\ e\ d,\ n\ n}\rightarrow{\color[rgb]{1,0,0}nd}$       $d^{2}+n^{2}$
      KRP       ${\color[rgb]{0,0,1}d},\mathbf{n\ d}\rightarrow{\color[rgb]{1,0,0}nd}$       $nd$
      KP       ${\color[rgb]{0,0,1}d},\mathbf{d\ d,\ n}\rightarrow{\color[rgb]{1,0,0}nd}$       $d^{2}+n$
@@
After obtaining the expanded $\mathbf{g},\mathbf{i},\mathbf{o}$, we feed them into element-wise gated linear recurrent layers as in Eq. 1 and Eq. 2, resulting in the output vector $\mathbf{y}_{t}\in\mathbb{R}^{n\times d}$. To project the expanded dimension back to the original dimension, we simply sum over the dimension corresponding to $n$.
@@
The results are shown in Table 3. We found that state expansion generally improves performance, with the low-rank matrix performing the best among these candidates.
Table 3: PESE Ablation. Ablation studies on various parameter-efficient methods, as described in Table 2. Each model was trained on 10 billion tokens from the Pile dataset.
Method n PPL Params (M)
Xfmr - 5.16 380
Xfmr++ - 4.62 386
HGRN1 1 5.10 379
LR 4 4.76 385
8 4.77 386
GLT 4 5.06 386
GLTI 4 4.83 386
KRP 4 5.08 386
KP 4 5.06 386
HGRN2 4 4.79 385
8 4.73 385
128 4.62 385
@@
However, these methods face training inefficiency issues, as they require conducting element-wise linear recurrence in high dimensions (i.e., $nd$). Since these element-wise operations cannot leverage tensor cores (a fast matrix multiplication unit on GPUs), the dramatically increasing FLOPs and I/O costs significantly slow down training when $n$ is large. We notice that this is similar to the case in Mamba [^1] Though Mamba has an attention mechanism (Ali et al., 2024) similar to that in linear attention, the attention computation cannot be expressed as a matrix multiplication like linear attention, and thus does not facilitate tensor core-based GPU acceleration, as well acknowledged in Mamba2 (Dao & Gu, 2024). , which requires a relatively small expansion ratio (i.e., $n=16$) and a custom I/O-efficient CUDA implementation to achieve a reasonably fast running speed.
@@
In the next subsection, we explore an alternative strategy that does not replace the dense projection matrices with structured ones but instead changes the element-wise gating operations in Eq.1 to other matrix/vector operations similar to those used in linear attention. This approach allows for more efficient training.
@@ 3.2 HGRN2
Figure 1: Network Structure of HGRN2. Each HGRN2 layer includes a token mixer layer, HGRU2, and a channel mixer, GLU. HGRU2 employs recurrent computation as described in Eq. 3, where $\mathbf{i}_{t}$ is the input state, $\mathbf{g}_{t}$ is the forget gate, $\mathbf{o}_{t}$ is the output gate, and $\beta^{i}$ is the lower bound for layer $i$.
The modification from HGRN1 to HGRN2 is simple yet effective. For the input gate, HGRN2 replaces the element-wise product with the outer product for state expansion. Consequently, $\mathbf{h}_{t}\in\mathbb{R}^{d\times d}$, and HGRN2 first diagonalizes the forget gate vector and uses the matrix dot product to update the hidden state. For the output gate, HGRN2 replaces the element-wise product with matrix-vector multiplication to project the expanded state back to the original dimension. The recurrent equation of HGRN2 is as follows:
$\mathbf{h}_{t}$ $=\mathbf{h}_{t-1}\cdot\mathrm{Diag}\{\mathbf{f}_{t}\}+\mathbf{i}_{t}\otimes(1-\mathbf{f}_{t})\in\mathbb{R}^{d\times d},$ (3)
$\mathbf{y}_{t}$ $=\mathbf{h}_{t}\cdot\mathbf{o}_{t}\in\mathbb{R}^{d},$
where $\mathrm{Diag}$ denotes the diagonalization of vectors, $\cdot$ represents the matrix dot product, and $\otimes$ indicates the outer product.
@@
The complexity of recurrence increases dramatically from $O(BNd)$ to $O(BNd^{2})$ due to state expansion. To address this, we introduce a multihead variant of HGRN (similar to that in linear attention) such that the complexity is reduced to $O(BNd^{2}/H)$ for the number of heads $H$, effectively making the state size $d^{2}/H$, i.e., the expansion ratio $n=d_{h}=d/H$ [^2] See Bolya et al. (2022) for more detailed complexity analysis. We conducted an ablation study on the expansion ratio (or head dimension) $n=\frac{d}{H}$, as shown in Figure 2. The results show that state expansion significantly improves language modeling performance. However, when the head dimension (i.e., state expansion ratio) exceeds 128, the performance gain diminishes. To balance computational cost and performance, we chose $d_{h}=128$ for the main experiments.
Figure 2: Expand Ratio (Head Dimension) Ablation. We tested the relationship between PPL (Perplexity) and the expand ratio on the Wikitext-103 (Merity et al., 2017) dataset (left) and a subset of the Pile (Gao et al., 2020) dataset (right).
@@
It is important to note that the recurrence form in HGRN2 is identical to that of GLA (Yang et al., 2023), except for the specific parameterization. We list the correspondences between the two parameterizations in Table 4. As shown, the output gate in HGRN2 corresponds to the query in GLA, while the output gate in GLA is omitted in HGRN2. The key vector in GLA corresponds to the input gate in HGRN2, which is tied to the forget gate, thereby saving parameters.
Table 4: The correspondence between HGRN2 and GLA is as follows.
@@
                 --                  $\mathbf{o}$ (output gate)
@@
Due to its computational structure’s similarity to GLA, we can directly leverage their chunkwise algorithm and highly optimized kernels for hardware-efficient large-scale training. For more details, we refer readers to their paper.
Concluding Remarks.
Although HGRN2 shares many similarities with GLA, we believe that HGRN2 offers a unique perspective distinct from linear attention, originating from the approach of gated linear RNNs. For instance, it may not be immediately clear from the perspective of linear attention why key vectors should be constrained within the range of (0, 1) or why the key vector and forget gate value should sum to one. However, these concepts become quite intuitive when starting from the gated linear RNN framework and exploring state expansion.